# 22footnotetext: Corresponding author

### v1

@@ 2.1. State Space Models
2.2. Structured State Space Sequence Models (S4)
S4 proposed by Gu et al. (Gu et al., 2021a) demonstrates how to efficiently compute all forms of the SSM: the recurrent representation, and the convolutional representation. Additionally, S4 employs a bilinear method for discretizing the parameters of the SSM, converting the state space parameter A into an approximation $\overline{\textbf{{A}}}$. In S4, the parameter $D$ from the original state space model is either set to 0 or used as a residual connection.
The discrete SSM can be expressed in its recurrent form as shown in Equation 3 and Equation 4.
(3) $x_{k}=\overline{\textbf{{A}}}x_{k-1}+\overline{\textbf{{B}}}u_{k}\;\;\;\;\;\overline{\textbf{{A}}}=(\mathbf{I}-\Delta/2\cdot\textbf{{A}})^{-1}(\mathbf{I}+\Delta/2\cdot\textbf{{A}})$
(4) $y_{k}=\overline{\textbf{{C}}}x_{k}\;\;\;\;\;\;\;\overline{\textbf{{B}}}=(\mathbf{I}-\Delta/2\cdot\textbf{{{A}}})^{-1}(\Delta\textbf{{B}})\;\;\;\;\;\;\;\overline{\textbf{{C}}}=\textbf{{C}}$
(5) $x_{0}=\overline{\textbf{{B}}}u_{0}\;\;\;\;\;\;x_{1}=\overline{\textbf{{AB}}}u_{0}+\overline{\textbf{{B}}}u_{1}\;\;\;\;\;\;x_{2}=\overline{\textbf{{A}}}^{2}\overline{\textbf{{B}}}u_{0}+\overline{\textbf{{AB}}}u_{1}+\overline{\textbf{{B}}}u_{2}$
(6) $y_{0}=\overline{\textbf{{CB}}}u_{0}\;\;\;\;\;\;y_{1}=\overline{\textbf{{CAB}}}u_{0}+\overline{\textbf{{CB}}}u_{1}\;\;\;\;\;\;y_{2}=\overline{\textbf{{CA}}}^{2}\overline{\textbf{{B}}}u_{0}+\overline{\textbf{{CAB}}}u_{1}+\overline{\textbf{{CB}}}u_{2}$
(7) $y_{k}=\overline{\textbf{{CA}}}^{k}\overline{\textbf{{B}}}u_{0}+\overline{\textbf{{CA}}}^{k-1}\overline{\textbf{{B}}}u_{1}+......+\overline{\textbf{{CAB}}}u_{k-1}+\overline{\textbf{{CB}}}u_{k}$
(8) $y=\overline{\textbf{{K}}}*u.$
(9) $\overline{\textbf{{K}}}\;\in\;\mathbb{R}^{L}:=\kappa_{L}(\overline{\textbf{{A}}},\overline{\textbf{{B}}},\overline{\textbf{{C}}}):=(\overline{\textbf{{CA}}}^{i}\overline{\textbf{{B}}})_{i\in[L]}=(\overline{\textbf{{CB}}},\overline{\textbf{{CAB}}},...,\overline{\textbf{{CA}}}^{L-1}\overline{\textbf{{B}}}).$
@@ 2.2. Structured State Space Sequence Models (S4)
S4 addresses the limitations of transformers by implementing these strategies. This empowers SSMs to excel in tasks requiring long-range dependencies such as Path-X (Tay et al., 2021) . In contrast, Transformers (et al., 2021; Katharopoulos et al., 2020b) in Path-X exhibit accuracy below 50% (worse than random guessing), whereas S4 achieves approximately 80% accuracy.
2.3. Simplified State Space Layers for Sequence Modeling (S5)
S5 proposed by Smith et al. (Smith et al., 2023), extends from S4 with similar initialization conditions, but enhances the architecture by adopting a Multiple Input Multiple Output (MIMO) approach. Notably, S5 introduces a learnable time scale parameter ($\Delta$), replacing the fixed parameter used in S4. Parameters in S5 are discretized using the Zero Order Hold (ZOH) method as mentioned in Equation 11, providing a refined parameter system compared to S4.
(11) $\overline{\mathbf{\Lambda}}=e^{\mathbf{\Lambda}\Delta},\;\;\;\;\;\overline{\textbf{{B}}}=\mathbf{\Lambda}^{-1}(\overline{\mathbf{\Lambda}}-\mathbf{I})\tilde{\textbf{{B}}},\;\;\;\;\;\overline{\textbf{{C}}}=\tilde{\textbf{{C}}},\;\;\;\;\;\overline{\textbf{{D}}}=\tilde{\textbf{{D}}}.$
In terms of computation, S5 employs a fully recurrent connection with parallel scanning , as detailed in Equation 3 and Equation 4. The authors highlight that a smaller latent space employs SSM to do parallel scanning where an associative operation is used in between during offline setting. This characteristic positions S5 for both online and offline processing, highlighting its utility in recurrent tasks within the time domain.
2.4. Selective Structured State Space Models (S6)
Extended from S5, S6 introduced by Gu et al. (Gu and Dao, 2024) incorporates SSMs within the Mamba Architecture. S6 builds upon the foundational assumptions of previous SSMs by leveraging a projection of the input function (using a linear layer) for initializing parameters $B$ and $C$. Notably, S6 also applies this projection to the step size parameter i.e., $\Delta$. To enhance computational efficiency, S6 implements faster recurrence connections using operations on the Static Random Access Memory (SRAM) of the GPU, with storage on the High Bandwidth Memory (HBM) similar to the principles outlined in FlashAttention mechanism by Dao et al. (Dao et al., 2022) and FlashAttention-2 mechanism by Dao et al. (Dao, 2023). The parameters in S6 are discretized using the Zero Order Hold method, following the approach set by S5. The Mamba architecture integrates components from both H3(Fu et al., 2023) and Gated MLP, incorporating an additional SSM layer and connections resembling the green parallelograms found in H3. Algorithm 1 & Algorithm 2 outlined below shows the differences between S4 and S6.
Algorithm 1 SSM (S4)
Input :$\textbf{Input :}\;\;x\;:\;(B,L,D)$
Output :$\textbf{Output :}\;\;y\;:\;(B,L,D)$
1:$\textbf{{A}}:(D,N)\leftarrow\text{Parameter}$triangleright Represents structured $N\times N$ matrix
2:$\textbf{{B}}:(D,N)\leftarrow\text{Parameter}$
3:$\textbf{{C}}:(D,N)\leftarrow\text{Parameter}$
4:$\mathbf{\Delta}:(D)\leftarrow\mathbf{\tau}_{\Delta}(\text{Parameter})$
5:$\overline{\textbf{{A}}},\overline{\textbf{{B}}}:(D,N)\leftarrow$discretize$(\Delta,\textbf{{A}},\textbf{{B}})$
6:$\textbf{{y}}\leftarrow SSM(\overline{\textbf{{A}}},\overline{\textbf{{B}}},\textbf{{C}})(x)$$\triangleright$ Time-invariant: recurrence or convolution
@@ 2.4. Selective Structured State Space Models (S6)
Input :$\textbf{Input :}\;\;x\;:\;(B,L,D)$
Output :$\textbf{Output :}\;\;y\;:\;(B,L,D)$
1:$\textbf{{A}}:(D,N)\leftarrow\text{Parameter}$triangleright Represents structured $N\times N$ matrix
2:B:($\textbf{{B}}:(B,L,N)\leftarrow s_{\mathbf{B}}(x)$
3:$\textbf{{C}}:(B,L,N)\leftarrow s_{\mathbf{C}}(x)$
4:$\mathbf{\Delta}:(B,L,D)\leftarrow\mathbf{\tau}_{\Delta}(\text{Parameter}+s_{\mathbf{\Delta}}(x))$
5:$\overline{\textbf{{A}}},\overline{\textbf{{B}}}:(B,L,D,N)\leftarrow$discretize$(\Delta,\textbf{{A}},\textbf{{B}})$
6:$\textbf{{y}}\leftarrow SSM(\overline{\textbf{{A}}},\overline{\textbf{{B}}},\textbf{{C}})(x)$$\triangleright$ Time-varying: recurrence (scan) only
@@
Mamba integrates a selective mechanism into its state space models (S6) to prioritize important content dynamically during training. SSM + Selection (S6) outlines a computational process for handling input sequences $x$ and generating corresponding output sequences $y$. It begins by initializing a structured matrix A and projecting the input sequence $x$ into tensors B and C using specific functions $s_{\mathbf{B}}$ and $s_{\mathbf{C}}$. The parameter $\mathbf{\Delta}$, serves as a time step parameter which is used in discretization. It is computed based on a function $\mathbf{\tau}$ incorporating additional parameters and $s_{\mathbf{\Delta}}$. Subsequently, A and B undergo discretization alongside $\mathbf{\Delta}$, resulting in transformed tensors $\overline{\textbf{{A}}}$ and $\overline{\textbf{{B}}}$ as mentioned below in Equation 12.
(12) $\overline{\mathbf{A}}=exp(\Delta\mathbf{A})\;\;\;\;\;\;\;\;\;\;\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}$
@@
Figure 4. Architecture of Mamba Block (Gu and Dao, 2024) combining H3 (Fu et al., 2023) and Gated MLP
3. Medical Image Analysis using Mamba
In this section, we cover the categorization of literature related to Mamba architectures, explore the optimizations that enhance their performance, and discuss various techniques and adaptations that expand their capabilities. Furthermore, we examine scanning techniques pertinent to Mamba and conclude by showcasing its diverse and impactful applications in the medical field.
3.1. Mamba Architectures
@@ 3.1.1. Pure Mamba
Vision Mamba (ViM) proposed by Zhu et al. (Zhu et al., 2024) incorporates bidirectional SSM by combining convolution with S6 in both forward and backward direction. Moreover, softplus (Zhao et al., 2018) function is applied over the selective scan parameter $\Delta_{0}$ to make sure the parameter stays positive. The parameters of the S6 are discretized with $\Delta_{0}$ parameter mentioned above. ViM employs a training strategy similar to Vision Transformers (ViT) (Dosovitskiy et al., 2021) where patches of inputs are tokenized by separating them into non-overlapping patches and applying a convolution layer on each patch with dimension $d$. The tokenized patches are then concatenated with class labels and learnable position encoding are added to the class label and tokenzied patches. Concatenating the class label in the middle has proven to work well as it exploits the recurrent nature of networks. Overall, ViM shows comparable differences in memory and performance to parametric heavy models such as DeiT-Ti (Data-efficient image Transformers-Tiny), DeiT-S (Data-efficient image Transformers-Small) proposed by Touvron et al. (Touvron et al., 2021) on tasks such as object detection, classification and segmentation. Particularly on higher dimensional images such as $1248\times 1248$, ViM consumes 73.2% less memory than DeiT and 2.8×2.8$\times$2.8 × faster than DeiT. Figure LABEL:fig:viM depicts the architecture of the ViM block, illustrating its key elements and their role in improving the model’s performance in vision tasks.
@@
Figure 6. Architectures of LABEL:fig:viM Vision Mamba (ViM) (Zhu et al., 2024) and LABEL:fig:vss Visual State Space (VSS) Block (Liu et al., 2024a)
@@
VMamba proposed by Liu et al. (Liu et al., 2024a) employs an architecture similar to transformers, replacing the traditional multihead attention block (Vaswani et al., 2017) with a novel approach. It utilizes a Visual State Space (VSS) block, which differentiates it from the standard Mamba architecture by incorporating depthwise convolution and SS2D (2D selective scan). This new implementation features a non-multiplicative branching method and replaces S6 used in Mamba with SS2D. While S6 performs well for NLP tasks, extending it to 2D vision data presents challenges in making S6 modules scan-independent. SS2D addresses this by implementing a gating mechanism, eliminating the need for branched multiplication as proposed in Mamba and ViM. In VMamba, patches are initially partitioned within the stem module, resulting in a feature map of size $H/4\times W/4$. As the data progresses through the layers, the feature map dimensions change sequentially to $H/8\times W/8$, $H/16\times W/16$, and finally $H/32\times W/32$, with C representing the network’s arbitrary dimensionality. Each stage, except the first, includes a downsampling block alongside VSS. Ultimately, a prediction head is employed to generate outputs for the designated task. SS2D consists of two stages: 1) cross scan module 2) cross merge module. The cross scan module in VMamba scans patches in four different directions: top to bottom, bottom to top, left to right and right to left. For each direction, an independent SSM is utilized, and the representations from these SSMs are combined using cross merging. On the performance perspective, VMamba achieves superior metrics with a minimal number of parameters and memory usage. Its performance is comparable to ViM-S, DeiT-S, and DeiT-B across various tasks including semantic segmentation, classification, and object detection. Figure LABEL:fig:vss depicts the architecture of the Visual State Space (VSS) block, showcasing its key components like SS2D scanning and depthwise convolution.
@@
Plain Mamba proposed by Yang et al. (Yang et al., 2024a) is a non-hierarchical SSM, similar to Vision Transformers (ViT). Drawing inspiration from ViT, Plain Mamba begins with patch embedding combined with position embedding, followed by a plain Mamba layer. In contrast to VSS, the plain Mamba layer utilizes gated multiplication of features, similar to ViM, but instead of a single SSM block, it employs four SSMs with continuous 2D scanning. The scan orders are the same as in VMamba, ensuring no positional bias and promoting uniform image understanding. Plain Mamba as shown in Figure LABEL:fig:plain, also incorporates direction-aware updating by embedding 2D relative position information into a flattened 1D layer within the SSM. Unlike ViM, it uses global pooling, followed by a classification head on top.
@@
Figure 8. Architectures of LABEL:fig:plain Plain Mamba (Yang et al., 2024a) and LABEL:fig:evss Efficient Visual State Space (EVSS) Block (Pei et al., 2024)
@@
Extension of VMamba: Pei et al. (Pei et al., 2024) proposed the EVSS block, which stands for Efficient Visual State Space, combines local and global features effectively. EVSS block shown in Figure LABEL:fig:evss, employs the ES2D module to extract the global feature map and a depth-wise convolutional branch to obtain the local feature map. It also incorporates a squeeze excitation block similar to SqueezeNet which is proposed by Iandola et al. (Iandola et al., 2016). The global and local features are then combined. ES2D introduces a novel method of patch-wise scanning. Initially, the input image is divided into patches belonging to different groups. These patches undergo forward and backward 2D scanning, similar to the VSS method, and are then passed to S6 and merged. In the overall architecture, EVSS blocks are utilized in the initial stages, while inverted residual blocks are employed in later stages. In conclusion, EfficientVMamba achieves better performance with less number of parameters compared to transformer-based and convolution-based models.
Figure 9. Architecture of Mamba ND Block (Li et al., 2024c)
@@
Li et al. (Li et al., 2024c) introduced the Mamba ND Block, which evaluates different scanning methods for 2D and 3D image representations. They found that the most effective approach involves scanning forward and then backward across height, width, and volume/time (if applicable). Separate Mamba blocks are designated for each scanning direction. The Mamba ND models outperformed transformer-based models while utilizing fewer parameters, resulting in enhanced performance in tasks such as video classification. The architecture of Mamba ND block is detailed in Figure 9, showcasing its design and structural components.
@@
MHS-VM proposed by Ji et al. (Ji, 2024), introduced an organized approach to construct visual features within 2D image spaces using Multi-Head Scan (MHS) module. This module projects embeddings from the previous layer into multiple lower-dimensional subspaces, where selective scan is performed along distinct scan routes. The resulting sub-embeddings undergo integration and projection back into the high-dimensional space. Additionally, the module incorporates a Scan Route Attention (SRA) mechanism to improve its ability to discern complex structures. To validate its efficiency, the SS2D block replaces the original block in VM-UNet, resulting in significant performance improvements while reducing parameters. Figure LABEL:fig:local illustrates the architecture of MHS-VM, detailing its use of components such as depthwise convolution, layer normalization, and multihead scanning.
@@
Extension of ViM: Huang et al. (Huang et al., 2024a) proposed Local Mamba, which implements a novel scanning mechanism where tokens are partitioned into distinct windows, and scanning is performed continuously within each patch of these windows. Local Mamba consists of four SSM blocks with different scanning patterns: a $7\times 7$ local scan, a $2\times 2$ local scan, a vertical scan (both forward and backward), and a horizontal scan (both forward and backward). Figure LABEL:fig:mhsvm depicts the architecture of Local Mamba, which incorporates four SSM blocks.
Figure 11. Architectures of LABEL:fig:local Local Mamba Block (Huang et al., 2024a) and LABEL:fig:mhsvm MHS Block (Ji, 2024)
Mamba with ViM: Wang et al. (Wang and Ma, 2024) proposed Weak-Mamba-UNet, a hybrid model which integrates CNN, ViT, and ViM-based Mamba models to achieve medical image segmentation using scribble-based annotations. The model assigns specific weighting factors to each network, denoted by parameters $\alpha,\beta$ and $\gamma$ for respective models. T Mamba, proposed by Hao et al. (Hao et al., 2024), introduced a modified architecture based on ViM. This architecture includes an additional branched input representation, which undergoes a weighted fourier transformation and band-pass filtering, followed by an inverse fourier transformation. Using a gated selection unit, these transformed representations are then projected into a single unified representation.
Mamba with VSS: Wu et al. (Wu et al., 2024a) proposed H-vmunet, which presents a modified architecture based on the VSS block that enables selective scan 2D. The paper introduces Higher Order Visual State Space (H-VSS), employing both first order and third order operations. Each order is a mathematical multiplication of its own projection with a local-SS2D block. The local-SS2D block consists of a LayerNorm (Ba et al., 2016), followed by a $C/2$ convolutional layer and $C/2$ SS2D layer. These are concatenated, and another LayerNorm is applied at the end. Wu et al. (Wu et al., 2024b) proposed UltraLight VM-UNet which employs a parallel vision mamba layer. Initially, the input is normalized and split into the desired number of channel-wise parts. Each split part undergoes a Mamba layer with a residual connection using an adjustment factor. The results are then concatenated and projected out with LayerNorm. For skip connections, UltraLight VM-UNet uses a Spatial Attention Bridge and a Channel Attention Bridge, similar to the traditional approach. Light M-UNet proposed by Liao et al. (Liao et al., 2024) employs the RVM (Residual Vision Mamba) layer in the encoder block. This involves applying LayerNorm, followed by a VSS module, and then performing a scaled residual connection. The decoder block utilizes depthwise convolution with a scaled residual connection.
Others: SegMamba proposed by Xing et al. (Xing et al., 2024) has the TOM (Tri-orientated Spatial Mamba Block) module which improves the sequential modeling of 3D features from three different directions, after which the spatial features are modeled by a gated spatial convolution module. The model architecture includes a gated spatial convolution followed by LayerNorm and TOM block, which is residually connected with the input, creating an architecture similar to transformers. The Tri-Oriented Mamba block first flattens the 3D representation and applies Mamba layer to rows, columns, and features respectively. Following this, the encoder output is downsampled using Conv3D networks, and then a CNN-based decoder is utilized. Wang et al. (Wang et al., 2024a) proposed LKM-UNet where the image is partitioned into pixel-level segments and bidirectional Mamba is applied to each segment. After pooling the outputs, bidirectional Mamba is applied again in the network. Finally, the resulting representation is added residually to the pixel-level representation.
@@
A mechanism that plays a crucial role in Mamba architectures is scanning, which is detailed in the upcoming subsection 3.2.
3.1.2. Variants of U-Net
Figure 12. Architectural Comparison of Traditional U-Net and Mamba U-Net
The integration of Mamba blocks into the U-Net (Ronneberger et al., 2015) architecture has resulted in several variants which are designed to improve its performance. For instance, they can be added before the first encoder layer, after an encoder layer, within the skip connections or can even replace the whole encoder block in UNet architecture. Figure 12 presents a comparision of traditional U-Net and Mamba U-Net (Wang et al., 2024d) architecture. Both networks share a standard UNet architecture having encoder and decoder blocks connected by skip connections but Mamba U-Net incorporates mamba blocks.

### v3

@@ 2.4 Selective Structured State Space Models (S6)
[…]
@@
$\textbf{Input :}\;\;x\;:\;(B,L,D)$
$\textbf{Output :}\;\;y\;:\;(B,L,D)$
[…]
@@
The parameters in S6 are discretized using the Zero Order Hold method, following the approach set by S5. The Mamba architecture integrates components from both H3[32] and Gated MLP, incorporating an additional SSM layer and connections resembling the green parallelograms found in H3. Algorithm 1 & Algorithm 2 outlined above shows the differences between S4 and S6. Mamba integrates a selective mechanism into its state space models (S6) to prioritize important content dynamically during training. SSM + Selection (S6) outlines a computational process for handling input sequences $x$ and generating corresponding output sequences $y$. It begins by initializing a structured matrix A and projecting the input sequence $x$ into tensors B and C using specific functions $s_{\mathbf{B}}$ and $s_{\mathbf{C}}$. The parameter $\mathbf{\Delta}$, serves as a time step parameter which is used in discretization. It is computed based on a function $\mathbf{\tau}$ incorporating additional parameters and $s_{\mathbf{\Delta}}$. Subsequently, A and B undergo discretization alongside $\mathbf{\Delta}$, resulting in transformed tensors $\overline{\textbf{{A}}}$ and $\overline{\textbf{{B}}}$ as mentioned below in Equation 12.
$\overline{\mathbf{A}}=exp(\Delta\mathbf{A})\;\;\;\;\;\;\;\;\;\;\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}$ (12)
@@
Figure 5: Architecture of Mamba Block [15] combining H3 [32] and Gated MLP
3 Medical Image Analysis using Mamba
In this section, we cover the categorization of literature related to Mamba architectures, explore optimizations that enhance their performance, and discuss various techniques and adaptations that expand their capabilities. Furthermore, we examine scanning techniques pertinent to Mamba and conclude by showcasing its diverse and impactful applications in the medical field.
3.1 Mamba Architectures
@@ 3.1.1 Pure Mamba
Vision Mamba (ViM) proposed by Zhu et al. [97] incorporates bidirectional SSM by combining convolution with S6 in both forward and backward directions. Moreover, softplus [125] function is applied to the selective scan parameter $\Delta_{0}$ to make sure the parameter stays positive. The parameters of the S6 are discretized with the $\Delta_{0}$ parameter mentioned above. ViM employs a training strategy similar to Vision Transformers (ViT) [126] where patches of inputs are tokenized by separating them into non-overlapping patches and applying a convolution layer on each patch with dimension $d$. The tokenized patches are then concatenated with class labels, and learnable position encoding are added to the class label and tokenzied patches. Overall, ViM shows comparable differences in memory and performance to parametric heavy models such as DeiT-Ti (Data-efficient image Transformers-Tiny), DeiT-S (Data-efficient image Transformers-Small) proposed by Touvron et al. [127] on tasks such as object detection, classification, and segmentation. Particularly in higher-dimensional images such as $1248\times 1248$, ViM consumes 73.2% less memory than DeiT and $2.8\times$ faster than DeiT. Figure LABEL:fig:viM depicts the architecture of the ViM block, illustrating its key elements and their role in improving the model’s performance in vision tasks.
@@
Figure 7: Architectures of LABEL:fig:viM Vision Mamba (ViM) [97] and LABEL:fig:vss Visual State Space (VSS) Block [98]
@@
VMamba proposed by Liu et al. [98] employs an architecture similar to transformers, replacing the traditional multi-head attention block [12] with a novel approach. It utilizes a Visual State Space (VSS) block, which differentiates it from the standard Mamba architecture by incorporating depthwise convolution and SS2D (2D selective scan). This new implementation features a non-multiplicative branching method and replaces S6 used in Mamba with SS2D. While S6 performs well for NLP tasks, extending it to 2D vision data presents challenges in making S6 modules scan-independent. SS2D addresses this by implementing a gating mechanism, eliminating the need for branched multiplication as proposed in Mamba and ViM. In VMamba, patches are initially partitioned within the stem module, resulting in a feature map of size $H/4\times W/4$. As the data progresses through the layers, the feature map dimensions change sequentially to $H/8\times W/8$, $H/16\times W/16$, and finally $H/32\times W/32$, with C representing the network’s arbitrary dimensionality. Each stage, except the first, includes a downsampling block alongside VSS. Ultimately, a prediction head is employed to generate outputs for the designated task. SS2D consists of two stages: 1) cross scan module 2) cross merge module. The cross scan module in VMamba scans patches in four different directions: top to bottom, bottom to top, left to right and right to left. For each direction, an independent SSM is utilized, and the representations from these SSMs are combined using cross merging. On the performance perspective, VMamba achieves superior metrics with a minimal number of parameters and memory usage. Its performance is comparable to ViM-S, DeiT-S, and DeiT-B across various tasks, including semantic segmentation, classification, and object detection. Figure LABEL:fig:vss depicts the architecture of the Visual State Space (VSS) block, showcasing its key components such as SS2D scanning and depthwise convolution.
@@
Plain Mamba proposed by Yang et al. [102] is a non-hierarchical SSM, similar to Vision Transformers (ViT). Drawing inspiration from ViT, Plain Mamba begins with patch embedding combined with position embedding, followed by a plain Mamba layer. In contrast to VSS, the plain Mamba layer utilizes gated multiplication of features, similar to ViM, but instead of a single SSM block, it employs four SSMs with continuous 2D scanning. The scan orders are the same as in VMamba, ensuring no positional bias and promoting uniform image understanding. Plain Mamba as shown in Figure LABEL:fig:plain, also incorporates direction-aware updating by embedding 2D relative position information into a flattened 1D layer within the SSM. Unlike ViM, it uses global pooling, followed by a classification head on top.
@@
Figure 9: Architectures of LABEL:fig:plain Plain Mamba [102] and LABEL:fig:evss Efficient Visual State Space (EVSS) Block [101]
@@
Extension of VMamba: Pei et al. [101] proposed the EVSS block, which stands for Efficient Visual State Space, which combines local and global features effectively. EVSS block shown in Figure LABEL:fig:evss, employs the ES2D module to extract the global feature map and a depth-wise convolutional branch to obtain the local feature map. It also incorporates a squeeze excitation block similar to SqueezeNet which is proposed by Iandola et al. [128]. The global and local features are then combined. ES2D introduces a novel method of patch-wise scanning. Initially, the input image is divided into patches belonging to different groups. These patches undergo forward and backward 2D scanning, similar to the VSS method, and are then passed to S6 and merged. In the overall architecture, EVSS blocks are utilized in the initial stages, while inverted residual blocks are employed in the later stages. In conclusion, EfficientVMamba achieves better performance with less number of parameters compared to transformer-based and convolution-based models.
Figure 10: Architecture of Mamba ND Block [117]
@@
Li et al. [117] introduced the Mamba ND Block, which evaluates different scanning methods for 2D and 3D image representations. They found that the most effective approach involves scanning forward and then backward across height, width, and volume/time (if applicable). Separate Mamba blocks are designated for each scanning direction. The Mamba ND models outperformed transformer-based models while utilizing fewer parameters, resulting in enhanced performance in tasks such as video classification. The architecture of the Mamba ND block is detailed in Figure 10, showcasing its design and structural components.
@@
MHS-VM proposed by Ji et al. [105], introduced an organized approach to construct visual features within 2D image spaces using Multi-Head Scan (MHS) module. This module projects embeddings from the previous layer into multiple lower-dimensional subspaces, where selective scan is performed along distinct scan routes. The resulting sub-embeddings undergo integration and projection back into the high-dimensional space. Additionally, the module incorporates a Scan Route Attention (SRA) mechanism to improve its ability to discern complex structures. To validate its efficiency, the SS2D block replaces the original block in VM-UNet, resulting in significant performance improvements while reducing parameters. Figure LABEL:fig:local illustrates the architecture of MHS-VM, detailing its use of components such as depthwise convolution, layer normalization, and multihead scanning.
@@
Extension of ViM: Huang et al. [100] proposed Local Mamba, which implements a novel scanning mechanism where tokens are partitioned into distinct windows, and scanning is performed continuously within each patch of these windows. Local Mamba consists of four SSM blocks with different scanning patterns: a $7\times 7$ local scan, a $2\times 2$ local scan, a vertical scan (both forward and backward), and a horizontal scan (both forward and backward). Figure LABEL:fig:mhsvm depicts the architecture of Local Mamba, which incorporates four SSM blocks.
Figure 12: Architectures of LABEL:fig:local Local Mamba Block [100] and LABEL:fig:mhsvm MHS Block [105]
@@
A mechanism that plays a crucial role in Mamba architectures is scanning, which is detailed in the upcoming Section 3.2.
3.1.2 Variants of U-Net
Figure 13: Architectural Comparison of Traditional U-Net and Mamba U-Net